The atmospheric pressure due to water is 1.3 x 106 Nm-2. What is the total pressure at the bottom of the ocean 10m deep. (Density of water = 1000kgm-3, g=10ms-2).
1
Expert's answer
2020-05-08T15:47:44-0400

Solution.

"p_a=1.3\sdot10^6N/m^2;"

"h=10m;"

"\rho=1000kg/m^3;"

"g=10m\/s^2;"

The total pressure on the ocean floor is equal to the sum of atmospheric pressure and hydrostatic pressure.

"p=p_a+p_h;"

Hydrostatic pressure is sought by the formula:

"p_h=\rho gh;"

"p_h=1000kg/m^3\sdot10m/s^2\sdot10m=10^5N/m^2;"

"p=1.3\sdot10^6N/m^2+10^5N/m^2=1.4\sdot10^6N/m^2;"

Answer: "p=1.4\sdot10^6N/m^2."